Your refrigerator has a mass of 104.5 kg, including the food in it. It is standing in the middle of your kitchen, and you need to move it. The coefficients of static and kinetic friction between the fridge and the tile floor are 0.44 and 0.35, respectively. What is the magnitude of the force of friction acting on the fridge, if you push against it horizontally with a force of each magnitude? ( in number format with 4 decimal places)

(a) 300 N
(b) 500 N
(c) 700 N
Answer #1

here,


coefficient of staticc friction , us = 0.44

coefficient of kinetic friction , uk = 0.35

mass of the refridgerator , m = 104.5 kg


the magnitude of static frictional force , F = us * m * g

F = 0.44 * 104.5 * 9.8

F = 450.604 N

(a)

as the force applied , f = 300 N

therefore the friction force is 300 N

(b)

the magnitude of kinetic frictional force , F = uk * m * g

F = 0.35 * 104.5 * 9.8

F = 358.435 N

the magnitude of kinetic frictional force is 358.435 N

(c)

the magnitude of kinetic frictional force , F = uk * m * g

F = 0.35 * 104.5 * 9.8

F = 358.435 N

the magnitude of kinetic frictional force is 358.435 N